OpenStudy (anonymous):

2^3 X 2. (two to the third power times two) do i have to put a 1 as the exponent for the two without one?

OpenStudy (amistre64):

have to? no

OpenStudy (amistre64):

is it wrong to ^1 it? no

OpenStudy (amistre64):

its easier to work with sometimes if its there tho since b^n * b^m = b^n+m

OpenStudy (anonymous):

yes. exactly dontletmefade. so you get 2^3 * 2^1 = 2^4